User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!


Hobart 26m

Disk VSN: HAY-0050/6000/1024 | Data volume at beginning: 0.0 GB

  • 2200UT Cryo pressure is high, the temperature is high, dbbcifa has 0 attenuation since taken over of shift, Warren said not much we can do about the pressure, temperature gauge doesn't work, power is closed to target 40000 so it should be okay(Lim)
  • 0445UT clkoff maserdelay large offset, consistent drift started around 0310UT but mk5=dot? looked normal. Halt schedule for dbbc reset. (Lim)
  • 0450UT configured dbbc and the delay back to normal. bad scans probably from 350-0311b to 350-0448a(Lim)
You could leave a comment if you were logged in.
/home/www/auscope/opswiki/data/pages/handover/r1977.txt · Last modified: 2020/12/15 04:57 by Lim Chin Chuan